OpenStudy (anonymous):

write an equation of the line with given point and parallel to the given line write in y=mx+b form (-4,4) 9x=4y+3

OpenStudy (anonymous):

Parallel means the slopes are equal. So first, find the slope, m, of the given line. It happens to be m=(9/4). So far, if you write your line as y=mx+b, you have y=(9/4)x+b. Your point, (-4,4) can be your x and y. Substitute. y=mx+b 4=(9/4)(-4)+b 4=-9+b b=4+-9=-5. Your line has m=(9/4) and b=-5. Therefore, the line is y=(9/4)x-5.

OpenStudy (chaise):

9x = 4y +3 -4y = -9x +3 y = 2.25x + 0.75 this is the general equation to the line. If the line is parralel to the above line, then it will have the SAME gradient and a DIFFERENT y intercept. y = 2.25x + b 4 = 2.25x-4 + b 4 = -9 + b b = 13 y = 2.25x+13

OpenStudy (anonymous):

Chaise, not to be rude, but how did you go from 9x=4y+3 to -4y=-9x+3?